# Can Direct Latent Model Learning Solve Linear Quadratic Gaussian Control?

### v2

@@ 2.1 Latent model of LQG
where $b_{t}>0$ is a problem-dependent constant, and $\gamma_{t}=\|z^{\ast}_{t}-x_{t}\|_{Q^{\ast}_{t}}^{2}-b_{t}$, $\eta_{t}=\bigl{\langle}z^{\ast}_{t},x_{t}-z^{\ast}_{t}\bigr{\rangle}_{Q^{\ast}_{t}}$ are both zero-mean subexponential random variables. Under Assumptions 1 and 4, $b_{t}=\mathcal{O}(1)$ and $\|\gamma_{t}\|_{\psi_{1}}=\mathcal{O}(d_{x}^{1/2})$; moreover, if control $u_{t}\sim\mathcal{N}(0,\sigma_{u}^{2}I)$ for $0\leq t\leq T$, then $\|\eta_{t}\|_{\psi_{1}}=\mathcal{O}(d_{x}^{1/2})$.
@@
By the property of the Kalman filter, $z^{\ast}_{t}=\mathbb{E}[x_{t}\;|\;y_{0:t},u_{0:(t-1)}]$ is a function of the past history $(y_{0:t},u_{0:(t-1)})$. $u_{t}$ is a function of the past history $(y_{0:t},u_{0:(t-1)})$ and some independent random variables. Since $i_{t+1}$ is independent of the past history $(y_{0:t},u_{0:(t-1)})$, it is independent of $z^{\ast}_{t}$ and $u_{t}$. For the cost function,
$c_{t}=\|z^{\ast}_{t}\|_{Q^{\ast}_{t}}^{2}+\|u_{t}\|_{R^{\ast}_{t}}^{2}+\|z^{\ast}_{t}-x_{t}\|_{Q^{\ast}_{t}}^{2}+2\bigl{\langle}z^{\ast}_{t},x_{t}-z^{\ast}_{t}\bigr{\rangle}_{Q^{\ast}_{t}}.$
Let $b_{t}=\mathbb{E}[\|x_{t}-z^{\ast}_{t}\|_{Q^{\ast}_{t}}^{2}]$ be a constant that depends on system parameters $(A^{\ast}_{t},B^{\ast}_{t},\Sigma_{w_{t}})_{t=0}^{T-1}$, $(C^{\ast}_{t},\Sigma_{v_{t}})_{t=0}^{T}$ and $\Sigma_{0}$. Then, random variable $\gamma_{t}:=\|z^{\ast}_{t}-x_{t}\|_{Q^{\ast}_{t}}^{2}-b_{t}$ has zero mean. Since $(x_{t}-z^{\ast}_{t})$ is Gaussian, its squared norm is subexponential. Since $z^{\ast}_{t}$ and $(x_{t}-z^{\ast}_{t})$ are independent zero-mean Gaussian random vectors (Bertsekas, 2012), their inner product $\eta_{t}$ is a zero-mean subexponential random variable.
If the system is uniformly exponentially stable (Assumption 1) and the system parameters are regular (c.f. Assumption 4), then $(S^{\ast}_{t})_{t=0}^{T}$ given by RDE (2.3) has a bounded operator norm determined by system parameters $(A^{\ast}_{t},B^{\ast}_{t},C^{\ast}_{t},\Sigma_{w_{t}})_{t=0}^{T-1}$, $(\Sigma_{v_{t}})_{t=0}^{T}$ and $\Sigma_{0}$ (Zhang and Zhang, 2021). Since $S^{\ast}_{t}=\mathrm{Cov}(x_{t}-z^{\ast}_{t})$, $\|\gamma_{t}\|_{\psi_{1}}=\mathcal{O}(d_{x}^{1/2})$ by Lemma 7. By Assumption 1, if we apply zero control to the system, then $\|\mathrm{Cov}(z^{\ast}_{t})\|_{2}=\mathcal{O}(1)$. By Lemma 7, $\eta_{t}=\left\langle z^{\ast}_{t},x_{t}-z^{\ast}_{t}\right\rangle$ satisfies $\|\eta_{t}\|_{\psi_{1}}=\mathcal{O}(d_{x}^{1/2})$. ∎
@@
Now let us take a closer look at the state representation function. The Kalman filter can be written as $z^{\ast}_{t+1}=\overline{A}^{\ast}_{t}z^{\ast}_{t}+\overline{B}^{\ast}_{t}u_{t}+L^{\ast}_{t+1}y_{t+1}$, where $\overline{A}^{\ast}_{t}=(I-L^{\ast}_{t+1}C^{\ast}_{t+1})A^{\ast}_{t}$ and $\overline{B}^{\ast}_{t}=(I-L^{\ast}_{t+1}C^{\ast}_{t+1})B^{\ast}_{t}$. For $0\leq t\leq T$, unrolling the recursion gives
@@
Overall, a policy $\pi$ is a combination of state representation function $(M_{t})_{t=0}^{T-1}$ ($M_{T}$ is not needed) and feedback gain $(K_{t})_{t=0}^{T-1}$ in the latent model; in this case, we write $\pi=(M_{t},K_{t})_{t=0}^{T-1}$ as the composition of the two.
@@ 3 Methodology: cost-driven state representation learning
Our method falls into the cost-driven category, which is more direct than the other two types, in the sense that the cost is directly relevant to planning with a dynamic model, whereas the observation reconstruction functions and inverse models are not. Another reason why we call our method direct latent model learning is that compared with Markov parameter-based approaches for linear systems, our approach directly parameterizes the state representation function, without exploiting the structure of the Kalman filter, making our approach closer to empirical practice that was designed for general RL settings.
Reference (Subramanian et al., 2020) proposes to optimize a simple combination of cost and transition prediction errors to learn the approximate information state (Subramanian et al., 2020). That is, we parameterize a state representation function by matrices $(M_{t})_{t=0}^{T}$ and a latent model by matrices $((A_{t},B_{t})_{t=0}^{T-1},(Q_{t})_{t=0}^{T})$ and then solve
@@
for $10\leq t\leq T-$ 1 and $l_{T}^{(i)}=\big{(}\|M_{T}h_{T}^{(i)}\|_{Q_{T}}^{2}+b_{T}-c_{T}^{(i)}\big{)}^{2}$. The optimization problem (3.1) is nonconvex; even if we can find a global minimizer, it is unclear how to establish finite-sample guarantees for it. A main finding of this work is that for LQG, we can solve the cost and transition loss optimization problems sequentially, with the caveat of using cumulative costs.
Algorithm 1 Direct latent model learning
1:Input: sample size $n$, input noise magnitude $\sigma_{u}$, singular value threshold $\theta=\Theta(n^{-1/4})$
2:Collect $n$ trajectories using $u_{t}\sim\mathcal{N}(0,\sigma_{u}^{2}I)$, for $10\leq t\leq T-$ 1, to obtain data in the form of
@@
3:Run CoReL($\mathcal{D}_{\textnormal{raw}}$, $\theta$) (Algorithm 2) to obtain state representation function estimate $(\hat{M}_{t})_{t=0}^{T}$ and latent state estimates $(z_{t}^{(i)})_{t=0,i=1}^{T,n}$, so that the data are converted to
$\mathcal{D}_{\textnormal{state}}=(z_{0}^{(i)},u_{0}^{(i)},c_{0}^{(i)},\ldots,z_{T-1}^{(i)},u_{T-1}^{(i)},c_{T-1}^{(i)},z_{T}^{(i)},c_{T}^{(i)})_{i=1}^{n}$
4:Run SysId($\mathcal{D}_{\textnormal{state}}$) (Algorithm 3) to obtain system parameter estimates $((\hat{A}_{t},\hat{B}_{t})_{t=0}^{T-1},(\hat{Q}_{t})_{t=0}^{T})$
@@
Our method is summarized in Algorithm 1. It has three steps: cost-driven state representation function learning (CoReL, Algorithm 2), latent system identification (SysId, Algorithm 3), and planning by RDE (2.4).
@@
CoReL (Algorithm 2) is the core of our algorithm. Once the state representation function $(\hat{M}_{t})_{t=0}^{T}$ is obtained, SysId (Algorithm 3) identifies the latent system using linear and quadratic regression, followed by planning using RDE (2.4) to obtain controller $(\hat{K}_{t})_{t=0}^{T-1}$ from $((\hat{A}_{t},\hat{B}_{t},R^{\ast}_{t})_{t=0}^{T-1},$ $(\hat{Q}_{t})_{t=0}^{T})$. SysId consists of the standard regression procedures; the full algorithmic detail is shown in Algorithm 3. Below we explain the cost-driven state representation learning algorithm (CoReL, Algorithm 2) in detail.
@@ 3.1 Learning the state representation function
Algorithm 2 CoReL: cost driven state representation learning
1:Input: raw data $\mathcal{D}_{\textnormal{raw}}$, singular value threshold $\theta=\Theta((\ell(d_{y}+d_{u}))^{1/2}d_{x}^{3/4}n^{-1/4})$
@@
where $k=1$ for $10\leq t\leq l-$ 1 and $k=m\wedge(T-t+1)$ for $\ell\leq t\leq T$
3:Find $\widetilde{M}_{t}\in\mathbb{R}^{d_{x}\times((t+1)d_{y}+td_{u})}$ such that $\widetilde{M}_{t}^{\top}\widetilde{M}_{t}$ is an approximation of $\hat{N}_{t}$
4:For all $10\leq t\leq\ell-$ 1, set $\hat{M}_{t}=\textsc{TruncSV}{}(\widetilde{M}_{t},\theta)$; for all $\ell\leq t\leq T$, set $\hat{M}_{t}=\widetilde{M}_{t}$
@@
Algorithm 3 SysId: system identification
1:Input: data in the form of $(z_{0}^{(i)},u_{0}^{(i)},c_{0}^{(i)},\ldots,z_{T-1}^{(i)},u_{T-1}^{(i)},c_{T-1}^{(i)},z_{T}^{(i)},c_{T}^{(i)})_{i=1}^{n}$
2:Estimate the system dynamics by $(\hat{A}_{t},\hat{B}_{t})_{t=0}^{T-1}\in$triangleright pick min. Frobenius norm solution by pseudoinverse
$\operatorname*{argmin}_{(A_{t},B_{t})_{t=0}^{T-1}}\sum_{t=0}^{T-1}\sum_{i=1}^{n}\|A_{t}z_{t}^{(i)}+B_{t}u_{t}^{(i)}-z_{t+1}^{(i)}\|^{2}$ (3.4)
3:For all $10\leq t\leq\ell-$ 1 and $t=T$, set $\hat{Q}_{t}=I_{d_{x}}$
@@
The state representation function is learned via CoReL (Algorithm 2). Given the raw data consisting of $n$ trajectories, CoReL first solves the regression problem (3.3) to recover the symmetric matrix $\hat{N}_{t}$. The target $\overline{c}_{t}$ of regression (3.3) is defined by
@@
where $k=1$ for $10\leq t\leq\ell-$ 1 and $k=m\wedge(T-t+1)$ for $\ell\leq t\leq T$. The superscript in $\overline{c}_{t}^{(i)}$ denotes the observed $\overline{c}_{t}$ in the $i$th trajectory. The quadratic regression has a closed-form solution, by converting it to linear regression using $\|v\|_{P}^{2}=\left\langle vv^{\top},P\right\rangle_{F}=\left\langle\mathrm{svec}(vv^{\top}),\mathrm{svec}(P)\right\rangle$.
@@
It is easy to verify that under the normalized parameterization the system satisfies Assumptions 123, and 4, up to a change of some constants in the bounds. Without loss of generality, we assume system (2.1) is in the normalized parameterization; if not, the recovered state representation function and latent system are with respect to the normalized parameterization.
Low-rank approximate factorization. Regression (3.3) has a closed-form solution; solving it gives $(\hat{N}_{t},\hat{b}_{t})_{t=0}^{T}$. Constants $(\hat{b}_{t})_{t=0}^{T}$ account for the variance of the state estimation error, and are not part of the state representation function; $d_{h}\times d_{h}$ symmetric matrices $(\hat{N}_{t})_{t=0}^{T}$ are estimates of $(M^{\ast}_{t})^{\top}M^{\ast}_{t}$ under the normalized parameterization, where $d_{h}=(t+1)d_{y}+td_{u}$. $M^{\ast}_{t}$ can only be recovered up to an orthogonal transformation, since for any orthogonal $S\in\mathbb{R}^{d_{x}\times d_{x}}$, $(SM^{\ast}_{t})^{\top}SM^{\ast}_{t}=(M^{\ast}_{t})^{\top}M^{\ast}_{t}$.
We want to recover $\widetilde{M}_{t}$ from $\hat{N}_{t}$ such that $\hat{N}_{t}=\widetilde{M}_{t}^{\top}\widetilde{M}_{t}$. Let $U\Lambda U^{\top}=\hat{N}_{t}$ be its eigenvalue decomposition. Let $\Sigma:=\max(\Lambda,0)$ be the positive semidefinite diagonal matrix containing nonnegative eigenvalues, where “$\max$” applies elementwise. If $d_{h}\leq d_{x}$, we can construct $\widetilde{M}_{t}=[\Sigma^{1/2}U^{\top};0_{(d_{x}-d_{h})\times d_{h}}]$ by padding zeros. If $d_{h}>d_{x}$, however, $\mathrm{rank}(\hat{N}_{t})$ may exceed $d_{x}$. Without loss of generality, assume that the diagonal elements of $\Sigma$ are in descending order. Let $\Sigma_{d_{x}}$ be the left-top $d_{x}\times d_{x}$ block of $\Sigma$ and $U_{d_{x}}$ be the left $d_{x}$ columns of $U$. By the Eckart-Young-Mirsky theorem, $\widetilde{M}_{t}=\Sigma_{d_{x}}^{1/2}U_{d_{x}}^{\top}$ provides the best approximation of $\hat{N}_{t}$ with $\widetilde{M}_{t}^{\top}\widetilde{M}_{t}$ among $d_{x}\times d_{h}$ matrices in terms of the Frobenius norm.
@@
=\;= $A^{\ast}_{t-1}(A^{\ast}_{t-2}z^{\ast}_{t-2}+B^{\ast}_{t-2}u_{t-2}+L^{\ast}_{t-1}i_{t-1})+B^{\ast}_{t-1}u_{t-1}+L^{\ast}_{t}i_{t}$
=\;= $[B^{\ast}_{t-1},\ldots,A^{\ast}_{t-1}A^{\ast}_{t-2}\cdots A^{\ast}_{t-\ell+1}B^{\ast}_{t-\ell}][u_{t-1};\ldots;u_{t-\ell}]+A^{\ast}_{t-1}A^{\ast}_{t-2}\cdots A^{\ast}_{t-\ell}z^{\ast}_{t-\ell}$
@@
where $(u_{\tau})_{\tau=t-\ell}^{t-1}$, $z^{\ast}_{t-\ell}$ and $(i_{\tau})_{\tau=t-\ell+1}^{t}$ are independent. The matrix multiplied by $[u_{t-1};\ldots;u_{t-\ell}]$ is precisely the controllability matrix $\Phi_{t-1,\ell}^{c}$. Then
$\mathrm{Cov}(z^{\ast}_{t})=\mathbb{E}[z^{\ast}_{t}(z^{\ast}_{t})^{\top}]\succcurlyeq\;$ $\Phi_{t-1,\ell}^{c}\mathbb{E}[[u_{t-1};\ldots;u_{t-\ell}][u_{t-1};\ldots;u_{t-\ell}]^{\top}](\Phi_{t-1,\ell}^{c})^{\top}$
=\;= $\sigma_{u}^{2}\Phi_{t-1,\ell}^{c}(\Phi_{t-1,\ell}^{c})^{\top}.$
@@
Since $h_{t}=[y_{0:t};u_{0:(t-1)}]$ and $(\mathrm{Cov}(y_{t}))_{t=0}^{T},(\mathrm{Cov}(u_{t}))_{t=0}^{T-1}$ have $\mathcal{O}(1)$ operator norms, by Lemma 8, $\|\mathrm{Cov}(h_{t})\|=\|\mathbb{E}[h_{t}h_{t}^{\top}]\|=\mathcal{O}(t)$. Hence,
@@
For $10\leq t\leq\ell-$ 1, $\sigma_{\min}^{+}(M^{\ast}_{t})\geq\beta>0$.
Still, for $10\leq t\leq\ell-$ 1, Assumption 5 does not guarantee the full-rankness of $\mathrm{Cov}(\widetilde{M}_{t}h_{t})$, not even a lower bound on its minimum positive singular value; that is why we introduce TruncSV that truncates the singular values of $\widetilde{M}_{t}$ by a threshold $\theta>0$. Concretely, we take $\hat{M}_{t}=(\mathbb{I}_{[\theta,+\infty)}(\Sigma_{d_{x}}^{1/2})\odot\Sigma_{d_{x}}^{1/2})U_{d_{x}}^{\top}$. Then, $\hat{M}_{t}$ has the same singular values as $\widetilde{M}_{t}$ except that those below $\theta$ are zeroed. We take $\theta=\Theta(\ell(d_{y}+d_{u})d_{x}^{3/4}n^{-1/4}\log^{1/4}(1/p))$ to ensure a sufficient lower bound on the minimum positive singular value of $\hat{M}_{t}$, without increasing the statistical errors.

### v3

@@ 2.1 Latent model of finite-horizon time-varying LQG
where $b_{t}>0$ is a problem-dependent constant, and $\gamma_{t}=\|x_{t}-z^{\ast}_{t}\|_{Q^{\ast}_{t}}^{2}-b_{t}$, $\eta_{t}=2\bigl{\langle}z^{\ast}_{t},x_{t}-z^{\ast}_{t}\bigr{\rangle}_{Q^{\ast}_{t}}$ are both zero-mean subexponential random variables. Under Assumptions 1 and 4, $b_{t}=\mathcal{O}(1)$ and $\|\gamma_{t}\|_{\psi_{1}}=\mathcal{O}(d_{x}^{1/2})$; moreover, if control $u_{t}\sim\mathcal{N}(0,\sigma_{u}^{2}I)$ for $0\leq t\leq T$, then $\|\eta_{t}\|_{\psi_{1}}=\mathcal{O}(d_{x}^{1/2})$.
@@
By the property of the Kalman filter, $z^{\ast}_{t}=\mathbb{E}[x_{t}\;|\;y_{0:t},u_{0:(t-1)}]$ is a function of the past history $(y_{0:t},u_{0:(t-1)})$. Action $u_{t}$ is a function of the past history $(y_{0:t},u_{0:(t-1)})$ and may contain noise independent of all other random variables. Innovation $i_{t+1}=C^{\ast}_{t+1}(A^{\ast}_{t}(x_{t}-z^{\ast}_{t})+w_{t})+v_{t+1}$ is independent of the past history $(y_{0:t},u_{0:(t-1)})$; hence, it is independent of $z^{\ast}_{t}$ and $u_{t}$. For the cost function,
$c_{t}=\|z^{\ast}_{t}\|_{Q^{\ast}_{t}}^{2}+\|u_{t}\|_{R^{\ast}_{t}}^{2}+\|x_{t}-z^{\ast}_{t}\|_{Q^{\ast}_{t}}^{2}+2\bigl{\langle}z^{\ast}_{t},x_{t}-z^{\ast}_{t}\bigr{\rangle}_{Q^{\ast}_{t}}.$
Let $b_{t}=\mathbb{E}[\|x_{t}-z^{\ast}_{t}\|_{Q^{\ast}_{t}}^{2}]$ be a constant that depends on system parameters $(A^{\ast}_{t},B^{\ast}_{t},\Sigma_{w_{t}})_{t=0}^{T-1}$, $(C^{\ast}_{t},\Sigma_{v_{t}})_{t=0}^{T}$ and $\Sigma_{0}$. Then, random variable $\gamma_{t}:=\|x_{t}-z^{\ast}_{t}\|_{Q^{\ast}_{t}}^{2}-b_{t}$ has zero mean. Since $(x_{t}-z^{\ast}_{t})$ is Gaussian, its squared norm is subexponential. Since $z^{\ast}_{t}$ and $(x_{t}-z^{\ast}_{t})$ are independent zero-mean Gaussian random vectors (Bertsekas, 2012), their inner product, and hence $\eta_{t}$, are zero-mean subexponential random variables.
If the system is uniformly exponentially stable (Assumption 1) and the system parameters are regular (Assumption 4), then $(S^{\ast}_{t})_{t=0}^{T}$ given by RDE (2.5) has a bounded operator norm determined by system parameters $(A^{\ast}_{t},B^{\ast}_{t},C^{\ast}_{t},\Sigma_{w_{t}})_{t=0}^{T-1}$, $(\Sigma_{v_{t}})_{t=0}^{T}$ and $\Sigma_{0}$ (Zhang and Zhang, 2021). Since $S^{\ast}_{t}=\mathrm{Cov}(x_{t}-z^{\ast}_{t})$, $\|\gamma_{t}\|_{\psi_{1}}=\mathcal{O}(d_{x}^{1/2})$ by Lemma 11. By Assumption 1, if we apply zero control to the system, then $\|\mathrm{Cov}(z^{\ast}_{t})\|_{2}=\mathcal{O}(1)$. By Lemma 11, $\eta_{t}=2\left\langle z^{\ast}_{t},x_{t}-z^{\ast}_{t}\right\rangle_{Q^{\ast}_{t}}$ satisfies $\|\eta_{t}\|_{\psi_{1}}=\mathcal{O}(d_{x}^{1/2})$. ∎
@@
Now let us take a closer look at the state representation function. The Kalman filter can be written as $z^{\ast}_{t+1}=\overline{A}^{\ast}_{t}z^{\ast}_{t}+\overline{B}^{\ast}_{t}u_{t}+L^{\ast}_{t+1}y_{t+1}$, where $\overline{A}^{\ast}_{t}=(I-L^{\ast}_{t+1}C^{\ast}_{t+1})A^{\ast}_{t}$ and $\overline{B}^{\ast}_{t}=(I-L^{\ast}_{t+1}C^{\ast}_{t+1})B^{\ast}_{t}$. For $0\leq t\leq T$, unrolling the recursion yields
@@
Overall, a policy $\pi$ is a combination of state representation function $(M_{t})_{t=0}^{T-1}$ ($M_{T}$ is not needed) and feedback gain $(K_{t})_{t=0}^{T-1}$ in the latent model, so we write $\pi=(M_{t},K_{t})_{t=0}^{T-1}$ as the composition of the two, and let $\pi^{\ast}=(M^{\ast}_{t},K^{\ast}_{t})_{t=0}^{T-1}$ denote the optimal policy.
@@ 3 Methodology: Cost-driven state representation learning
Our method falls into the cost-driven category. Compared with Markov parameter-based approaches for linear systems, our approach directly parameterizes the state representation function, without exploiting the structure of the Kalman filter, making our approach closer to empirical practice that was designed for general RL settings.
Subramanian et al. (2020) propose to optimize a simple combination of cost and transition prediction errors to learn the approximate information state (Subramanian et al., 2020). That is, we parameterize a state representation function by matrices $(M_{t})_{t=0}^{T}$ and a latent model by matrices $((A_{t},B_{t})_{t=0}^{T-1},(Q_{t})_{t=0}^{T})$ and then solve
@@
for $0\leq t\leq T-1$ and $l_{T}^{(i)}=\big{(}\|M_{T}h_{T}^{(i)}\|_{Q_{T}}^{2}+b_{T}-c_{T}^{(i)}\big{)}^{2}$. The optimization problem (3.1) is nonconvex; even if we can find a global minimizer, it is unclear how to establish finite-sample guarantees for it. A main finding of this work is that for LQG, we can solve the cost and transition loss optimization problems sequentially, with the caveat of using cumulative costs.
Algorithm 1 CoReL: cost-driven state representation learning
1:Input: sample size $n$, input noise magnitude $\sigma_{u}=\Theta(1)$, singular value threshold $\theta=\Theta(n^{-1/4})$ (hiding dependence on other problem parameters)
2:Collect $n$ trajectories using $u_{t}\sim\mathcal{N}(0,\sigma_{u}^{2}I)$, for $0\leq t\leq T-1$, to obtain data in the form of
@@
3:Run Algorithm 2 with $\mathcal{D}_{\textnormal{raw}}$ and $\theta$ to obtain state representation function estimate $(\hat{M}_{t})_{t=0}^{T}$ and latent state estimates $(\hat{z}_{t}^{(i)})_{t=0,i=1}^{T,n}$, so that the data are converted to
$\mathcal{D}_{\textnormal{state}}=(\hat{z}_{0}^{(i)},u_{0}^{(i)},c_{0}^{(i)},\ldots,\hat{z}_{T-1}^{(i)},u_{T-1}^{(i)},c_{T-1}^{(i)},\hat{z}_{T}^{(i)},c_{T}^{(i)})_{i=1}^{n}$
4:Run Algorithm 3 with $\mathcal{D}_{\textnormal{state}}$ to obtain system parameter estimates $((\hat{A}_{t},\hat{B}_{t})_{t=0}^{T-1},(\hat{Q}_{t})_{t=0}^{T})$
@@
Our method is summarized in CoReL (Algorithm 1). It has three steps: cost-driven state representation function learning (Algorithm 2), latent system identification (Algorithm 3), and planning by RDE (2.6).
@@
Algorithm 2 is the core of our algorithm. Once the state representation function $(\hat{M}_{t})_{t=0}^{T}$ is obtained, Algorithm 3 identifies the latent model using linear and quadratic regressions, followed by planning using RDE (2.6) to obtain the controller $(\hat{K}_{t})_{t=0}^{T-1}$ from $((\hat{A}_{t},\hat{B}_{t},R^{\ast}_{t})_{t=0}^{T-1},$ $(\hat{Q}_{t})_{t=0}^{T})$. Algorithm 3 consists of the standard regression procedures. We explain Algorithm 2 below.
@@ 3.1 Learning the state representation function
Algorithm 2 Cost-driven state representation function learning
1:Input: raw data $\mathcal{D}_{\textnormal{raw}}$, singular value threshold $\theta$
@@
where $k=1$ for $0\leq t\leq l-1$ and $k=m\wedge(T-t+1)$ for $\ell\leq t\leq T$
3:Find $\widetilde{M}_{t}\in\operatorname*{argmin}_{M\in\mathbb{R}^{d_{x}\times((t+1)d_{y}+td_{u})}}\|M^{\top}M-\hat{N}_{t}\|_{F}$
4:For all $0\leq t\leq\ell-1$, set $\hat{M}_{t}=\textsc{TruncSV}{}(\widetilde{M}_{t},\theta)$; for all $\ell\leq t\leq T$, set $\hat{M}_{t}=\widetilde{M}_{t}$
@@
Algorithm 3 Latent model identification
1:Input: data in the form of $(\hat{z}_{0}^{(i)},u_{0}^{(i)},c_{0}^{(i)},\ldots,\hat{z}_{T-1}^{(i)},u_{T-1}^{(i)},c_{T-1}^{(i)},\hat{z}_{T}^{(i)},c_{T}^{(i)})_{i=1}^{n}$
2:Estimate the system dynamics by $(\hat{A}_{t},\hat{B}_{t})_{t=0}^{T-1}\in$
$\operatorname*{argmin}_{(A_{t},B_{t})_{t=0}^{T-1}}\sum_{t=0}^{T-1}\sum_{i=1}^{n}\|A_{t}\hat{z}_{t}^{(i)}+B_{t}u_{t}^{(i)}-\hat{z}_{t+1}^{(i)}\|^{2},$ (3.4)
picking the minimal-Frobenius-norm solution by pseudoinverse, as in (4.6)
3:For all $0\leq t\leq\ell-1$ and $t=T$, set $\hat{Q}_{t}=I_{d_{x}}$
@@
The state representation function is learned via Algorithm 2. Given the raw data consisting of $n$ trajectories, Algorithm 2 first solves the regression problem (3.3) to recover the symmetric matrix $\hat{N}_{t}$. The target $\overline{c}_{t}$ of regression (3.3) is defined by
@@
where $k=1$ for $0\leq t\leq\ell-1$ and $k=m\wedge(T-t+1)$ for $\ell\leq t\leq T$. The superscript in $\overline{c}_{t}^{(i)}$ denotes the observed $\overline{c}_{t}$ in the $i$th trajectory. The quadratic regression has a closed-form solution, by converting it to linear regression using $\|v\|_{P}^{2}=\left\langle vv^{\top},P\right\rangle_{F}=\left\langle\mathrm{svec}(vv^{\top}),\mathrm{svec}(P)\right\rangle$.
@@
One can verify that under the normalized parameterization, the system satisfies Assumptions 123, and 4, up to a change of some constants in the bounds. Without loss of generality, we assume system (2.1) is in the normalized parameterization.
Low-rank approximate factorization. Regression (3.3) has a closed-form solution $(\hat{N}_{t},\hat{b}_{t})_{t=0}^{T}$. Constants $(\hat{b}_{t})_{t=0}^{T}$ account for the variance of the state estimation error, and are not part of the state representation function; $d_{h}\times d_{h}$ symmetric matrices $(\hat{N}_{t})_{t=0}^{T}$ are estimates of $(M^{\ast}_{t})^{\top}M^{\ast}_{t}$ under the normalized parameterization, where $d_{h}=(t+1)d_{y}+td_{u}$. $M^{\ast}_{t}$ can only be recovered up to an orthogonal transformation, as for any orthogonal $S\in\mathbb{R}^{d_{x}\times d_{x}}$, $(SM^{\ast}_{t})^{\top}SM^{\ast}_{t}=(M^{\ast}_{t})^{\top}M^{\ast}_{t}$.
We want to recover $\widetilde{M}_{t}$ from $\hat{N}_{t}$ such that $\hat{N}_{t}=\widetilde{M}_{t}^{\top}\widetilde{M}_{t}$. Let $U\Lambda U^{\top}=\hat{N}_{t}$ be its eigenvalue decomposition. Let $\Sigma:=\max(\Lambda,0)$ be the positive semidefinite diagonal matrix with nonnegative eigenvalues, where $\max$” applies elementwise. If $d_{h}\leq d_{x}$, we can construct $\widetilde{M}_{t}=[\Sigma^{1/2}U^{\top};0_{(d_{x}-d_{h})\times d_{h}}]$ by padding zeros. If $d_{h}>d_{x}$, however, $\mathrm{rank}(\hat{N}_{t})$ may exceed $d_{x}$. Without loss of generality, assume that the diagonal elements of $\Sigma$ are in descending order. Let $\Sigma_{d_{x}}$ be the top-left $d_{x}\times d_{x}$ block of $\Sigma$ and $U_{d_{x}}$ be the left $d_{x}$ columns of $U$. By the Eckart-Young-Mirsky theorem, $\widetilde{M}_{t}=\Sigma_{d_{x}}^{1/2}U_{d_{x}}^{\top}$ provides the best approximation of $\hat{N}_{t}$ with $\widetilde{M}_{t}^{\top}\widetilde{M}_{t}$ among $d_{x}\times d_{h}$ matrices in terms of the Frobenius norm distance.
@@
$=\;$ $A^{\ast}_{t-1}(A^{\ast}_{t-2}z^{\ast}_{t-2}+B^{\ast}_{t-2}u_{t-2}+L^{\ast}_{t-1}i_{t-1})+B^{\ast}_{t-1}u_{t-1}+L^{\ast}_{t}i_{t}$
$=\;$ $[B^{\ast}_{t-1},\ldots,A^{\ast}_{t-1}A^{\ast}_{t-2}\cdots A^{\ast}_{t-\ell+1}B^{\ast}_{t-\ell}][u_{t-1};\ldots;u_{t-\ell}]+A^{\ast}_{t-1}A^{\ast}_{t-2}\cdots A^{\ast}_{t-\ell}z^{\ast}_{t-\ell}$
@@
where $(u_{\tau})_{\tau=t-\ell}^{t-1}$, $z^{\ast}_{t-\ell}$ and $(i_{\tau})_{\tau=t-\ell+1}^{t}$ are independent. The matrix multiplied by $u_{t-1};\ldots;u_{t-\ell}]$ is precisely the controllability matrix $\Phi_{t-1,\ell}^{c}$. Then
$\mathrm{Cov}(z^{\ast}_{t})=\mathbb{E}[z^{\ast}_{t}(z^{\ast}_{t})^{\top}]\succcurlyeq\Phi_{t-1,\ell}^{c}\mathbb{E}[[u_{t-1};\ldots;u_{t-\ell}][u_{t-1};\ldots;u_{t-\ell}]^{\top}](\Phi_{t-1,\ell}^{c})^{\top}=\sigma_{u}^{2}\Phi_{t-1,\ell}^{c}(\Phi_{t-1,\ell}^{c})^{\top}.$
@@
Since $h_{t}=[y_{0:t};u_{0:(t-1)}]$ and $(\mathrm{Cov}(y_{t}))_{t=0}^{T},(\mathrm{Cov}(u_{t}))_{t=0}^{T-1}$ have $\mathcal{O}(1)$ operator norms by Lemma 12, $\|\mathrm{Cov}(h_{t})\|=\|\mathbb{E}[h_{t}h_{t}^{\top}]\|=\mathcal{O}(t)$. Hence,
@@
For $0\leq t\leq\ell-1$, $\sigma_{\min}^{+}(M^{\ast}_{t})\geq\beta>0$.
Still, for $0\leq t\leq\ell-1$, Assumption 5 does not guarantee the full-rankness of $\mathrm{Cov}(\widetilde{M}_{t}h_{t})$, not even a lower bound on its minimum positive singular value; that is why we introduce TruncSV that truncates the singular values of $\widetilde{M}_{t}$ by a threshold $\theta>0$. Concretely, we take $\hat{M}_{t}=(\mathbb{I}_{[\theta,+\infty)}(\Sigma_{d_{x}}^{1/2})\odot\Sigma_{d_{x}}^{1/2})U_{d_{x}}^{\top}$. Then, $\hat{M}_{t}$ has the same singular values as $\widetilde{M}_{t}$ except that those below $\theta$ are zeroed. We take $\theta=\Theta(\ell^{3/2}(d_{y}+d_{u})d_{x}^{3/4}n^{-1/4}\log^{1/4}(\ell/p))$ to ensure a sufficient lower bound on the minimum positive singular value of $\hat{M}_{t}$, without increasing the statistical errors.